OpenStudy (anonymous):

ill give medals Given equation A as 2x+1/4y=3 and equation B as 2/3x-y=6 , which expression will give the value of y? a-A − 3B b-A − 3/2 B c-2A − 3B d-1/3 A + B

OpenStudy (kc_kennylau):

Are there any missing parentheses?

OpenStudy (anonymous):

no

OpenStudy (kc_kennylau):

The equation of A is (2x+1)/4y=3, 2x+1/(4y)=3, or 2x+(1/4)y=3 ?